# Tight analyses of first-order methods with error feedback

### v1

@@ A Lyapunov function achieving the rate in (9), with $\xi^{\mathrm{EF}}$ defined in (7), is given by
Finally, the step size in (8) is worst-case optimal for $\mathrm{EF}$: $\forall\eta\geqslant 0$, we have $\rho^{\star}(\mathrm{EF}_{\eta})\geqslant\rho^{\star}(\mathrm{EF}_{\eta^{*}})$
Importantly, (9) shows that the rate is tight; that is, there exist $f\in\mathcal{F}_{\mu,L}$ and $(P,p)\in\mathbb{V}_{\ell}$ for which the rate is exactly achieved. Since the lower bound also applies to any other performance measure in our state space, our performance measure is optimal. This is formally demonstrated in the proof provided in Subsection B.1.
For completeness—and to support both our theoretical results and the tightness of the numerical results obtained through performance estimation—we provide additional figures comparing the empirically observed optimal step sizes for worst-case instances with our theoretical step size $\eta^{\star}$, across different values of $\epsilon$ and $\mu/L$. These results are shown in Figure 8, located in Subsection C.4. The numerical and analytical values match up to numerical accuracy.
@@ 3.3 Exact convergence rate and optimal tuning for $\mathrm{EF^{21}}$
We now state our main result on the $\mathrm{EF^{21}}$ algorithm, which is also tight within our class of candidate Lyapunov functions.
@@
Consider running Algorithm 3 with a compression operator satisfying Assumption 1 for some $\epsilon\in[0,1]$ on any function satisfying Assumptions 2, and 3. Let the step size be given by $\eta^{\star}$ in (8). Then, we have that
@@ A Lyapunov function achieving the rate in (12) is given by
$\mathcal{V}(\xi^{\mathrm{EF^{21}}},x;f)\coloneqq(1+\sqrt{\epsilon})\cdot\|g\|^{2}-2g^{\top}d+\|d\|^{2}=\|g-d\|^{2}+\sqrt{\epsilon}\cdot\|d\|^{2}.$ (13)
@@
The proof of this theorem is given in Appendix B.2. This second theoretical result analytically confirms the surprising numerical observation from Subsection 3.1, illustrated in Figure 1: $\mathrm{EF}$ and $\mathrm{EF^{21}}$ have the exact same optimal guarantee. Furthermore, the optimal step size is also the same, and can also be argued to be worst-case optimal in our setting using the same arguments as in Section 3.2. However, the optimal Lyapunov function is very different.
@@ 3.4 Tightness over multiple iterations and choice of state variables
A natural question arising from the above analysis concerns the tightness of the Lyapunov functions provided in Theorems 1 and 2 over multiple steps. Specifically, for $K\geqslant 2$, we investigate whether the convergence rate of a method $\mathcal{M}\in{\mathrm{EF},\mathrm{EF^{21}},\mathrm{CGD}}$ can be improved by analyzing $\mathcal{M}^{K}$, the method run over $K$ iterations, defined as follows:
@@
We provide a numerical answer to that question in Appendix C.2 for all algorithms discussed in the paper, showing that the analysis of the single-state Lyapunov functions used in this work are tight even if we consider multiple iterations. To that end, we plot the worst-case contractions for multiple iterations computed using PEPit [57], on the optimal Lyapunov functions given by (11) and (13).
@@ 4 Methodology
In this section, we present the methodology used to obtain the results in Section 3. The approach builds on the framework developed by [45]. We extend it to cover methods using compression operators under Assumption 1. We provide formal statements and proofs in Appendix A for completeness. Obtaining the proofs of Theorems 1 and 2 required a combination of advanced performance estimation techniques (finding optimal Lyapunov functions), several tricks, as well as symbolic computation and symbolic regression frameworks.
@@ 5 Conclusion
Our analysis is confined to the single-agent setting, both as a fundamental building block, and as a source of intuition for the multi-agent case. Importantly, $\mathrm{CGD}$ cannot serve as a baseline in the multi-agent setting as it fails to converge with more than one agent [24]. In contrast, $\mathrm{EF^{21}}$ was specifically designed to improve convergence over $\mathrm{EF}$ in the multi-agent setting. Yet, its convergence rate in the single-agent case matches exactly that of $\mathrm{EF}$—raising the question of how these two methods compare in a rigorous analysis with multiple agents. The findings of this paper raise two compelling questions:
@@ Appendix A Feasibility problems with compressors
This section presents the methodology used to search for Lyapunov functions. In a nutshell, we formulate the Lyapunov search problem as a quasi-convex optimization problem involving linear matrix inequalities. Those problems are typically solved through the use of an iterative procedure involving a binary search with semidefinite solvers—we use MOSEK [56] throughout. The main steps taken here can be viewed as a generalization of the procedure proposed in [45] to first-order methods using compression, in particular Algorithms 2 and 3. We also simplify a few steps that are not needed for our purposes. We start by reviewing the technique on a simpler example in Subsection A.1 before detailing the more tricky formulations involving compression.
@@ A.1 Feasibility problem for gradient descent
To introduce the concepts underlying the techniques used to construct Lyapunov functions for Algorithms 2 and 3, we begin with the simpler case of gradient descent on smooth, strongly convex functions. That is, we consider the algorithm:
@@
$0\geqslant\inf_{\lambda_{ij}\geqslant 0}$ 00 0 (25)
@@ A.2 Feasibility problem for $\mathrm{EF}$
Using the same ideas as in Subsection A.1, this section states the feasibility problem we solve to identify Lyapunov functions for Algorithm 2. In short, this requires adapting the two steps presented in Subsection A.1 to accommodate the compressed message. Practically speaking, Step 1 must be adapted to a slightly larger problem (with a few more states both in the Gram matrix $G$ and in the Lyapunov candidates to incorporate compression). Then, Step 2 follows directly by the same reasoning as before: the condition derived in Step 1 reduces to checking the feasibility of a linear matrix inequality that is linear in $(P,p)$, which in turn allows us to search for the Lyapunov function via binary search over $\rho$.
@@
Consider running Algorithm 2 with a compression operator satisfying Assumption 1 for some $\epsilon\in[0,1]$ on any function satisfying Assumptions 2 and 3. There exists a nonzero candidate Lyapunov function $\mathcal{V}_{(P,p)}$ of the form defined in (26), satisfying a given rate $\rho>0$, if and only if the following problem is feasible:
@@
The proof is decomposed into several steps that correspond to adapting the technical ingredients from Subsection A.1.
@@
$\Delta V_{P}(\rho)\coloneqq\begin{bmatrix}\bar{x}_{1}-\bar{x}_{\star}\\ \bar{g}_{1}-\bar{g}_{\star}\\ \bar{c}_{1}-\bar{c}_{\star}\\ \bar{e}_{1}-\bar{e}_{\star}\end{bmatrix}^{\top}P\begin{bmatrix}\bar{x}_{1}-\bar{x}_{\star}\\ \bar{g}_{1}-\bar{g}_{\star}\\ \bar{c}_{1}-\bar{c}_{\star}\\ \bar{e}_{1}-\bar{e}_{\star}\end{bmatrix}-\rho\begin{bmatrix}\bar{x}_{0}-\bar{x}_{\star}\\ \bar{g}_{0}-\bar{g}_{\star}\\ \bar{c}_{0}-\bar{c}_{\star}\\ \bar{e}_{0}-\bar{e}_{\star}\end{bmatrix}^{\top}P\begin{bmatrix}\bar{x}_{0}-\bar{x}_{\star}\\ \bar{g}_{0}-\bar{g}_{\star}\\ \bar{c}_{0}-\bar{c}_{\star}\\ \bar{e}_{0}-\bar{e}_{\star}\end{bmatrix},\quad$ (32)
@@
The interpolation conditions that enforce $f\in\mathcal{F}_{\mu,L}$ are identical to those we define in (23), but using the new basis vectors we define specifically for $\mathrm{EF}$. We do, however, need to introduce a new interpolation condition to encode the fact that we are using a contractive compressor. Using our basis vectors, this corresponds to introducing the matrices
@@
Finally, following the same reasoning as described in Step 1 and Step 2 of Subsection A.1, and using the technical modifications we outlined here, we arrive at the feasibility problem described in the statement of the lemma. ∎
@@ A.3 Feasibility problem for $\mathrm{EF^{21}}$
As in the previous section, one can now adapt the same ideas as in Subsection A.1 and Subsection A.2 to $\mathrm{EF^{21}}$. This section states the feasibility problem we solve to identify Lyapunov functions in this context.
@@
Lemma 2 ($\mathrm{EF}$ feasibility problem).
Consider running Algorithm 3 with a compression operator satisfying Assumption 1 for some $\epsilon\in[0,1]$ on any function satisfying Assumptions 2 and 3. There exists a nonzero candidate Lyapunov function $\mathcal{V}_{(P,p)}$ of the form defined in (37), satisfying a given rate $\rho>0$, if and only if the following problem is feasible:
@@
where matrices $(M_{ij})_{i,j\in\{0,1,\star\}}$ are defined as in Eq. 23, $(C_{i}^{\mathrm{EF^{21}}})_{i\in\{0,1\}}$ given below in (43), and $\Delta V_{P}(\rho),\Delta v_{p}(\rho)$ are given below in (44) and (45). Here $\geqslant_{2}$ denotes coordinate-wise inequality in $\mathbb{R}^{2}$.
@@
Finally, using Step 1 and Step 2 of Subsection A.1, we arrive at the feasibility problem described in the statement of the lemma. ∎
Appendix B Missing proofs

### v2

@@ A Lyapunov function achieving the rate in (9), with $\xi^{\mathrm{EF}}$ defined in (7), is given by
Finally, the step size in (8) is worst-case optimal for $\mathrm{EF}$: $\forall\eta\geqslant 0$, we have $\rho^{\star}(\mathrm{EF}_{\eta})\geqslant\rho^{\star}(\mathrm{EF}_{\eta^{*}})$ .
Importantly, (9) shows that the rate is tight; that is, there exist $f\in\mathcal{F}_{\mu,L}$ and $(P,p)\in\mathbb{V}_{\ell}$ for which the rate is exactly achieved. Since the lower bound also applies to any other performance measure in our state space, our performance measure is optimal. This is formally demonstrated in the proof provided in Subsection˜C.1. The proof was written using deterministic compressors, but the same guarantees hold under expectation with stochastic compressors as is explained in Subsection˜C.3.
For completeness—and to support both our theoretical results and the tightness of the numerical results obtained through performance estimation—we provide additional figures comparing the empirically observed optimal step sizes for worst-case instances with our theoretical step size $\eta^{\star}$, across different values of $\epsilon$ and $\mu/L$. These results are shown in Figure˜11, located in Subsection˜D.4. The numerical and analytical values match up to numerical accuracy.
@@ 3.3 Exact convergence rate and optimal tuning for $\mathrm{EF^{21}}$
We now state our main result on the $\mathrm{EF^{21}}$ algorithm, which is also tight.
@@
[Uncaptioned image][Uncaptioned image]
Consider running Algorithm 3 with a compression operator satisfying Assumption 1 for some $\epsilon\in[0,1]$ on any function satisfying Assumptions 2, and 3. Let the step size be given by $\eta^{\star}$ in (8). Then,
@@ A Lyapunov function achieving the rate in (12) is given by
$\mathcal{V}(\xi^{\mathrm{EF^{21}}},x;f)\coloneqq(1+\sqrt{\epsilon})\cdot\|g\|^{2}-2g^{\top}d+\|d\|^{2}=\|g-d\|^{2}+\sqrt{\epsilon}\cdot\|g\|^{2}.$ (13)
@@
Remark 1.
The explicit rate for $\mathrm{EF}$ and $\mathrm{EF^{21}}$ can be written as
$\rho=\sqrt{\epsilon}+\left(\frac{1-\sqrt{\epsilon}}{2}\right)\left(\frac{\kappa-1}{\kappa+1}\right)^{2}\left[1-\sqrt{\epsilon}+\sqrt{(1+\sqrt{\epsilon})^{2}+\sqrt{\epsilon}16\frac{\kappa}{(\kappa-1)^{2}}}\right],$ (14)
where $\kappa=L/\mu$. A detailed comparison between this and the existing rate for $\mathrm{EF^{21}}$ under the Łojasiewicz inequality [14] is provided in Appendix˜B.
The proof of this theorem is given in Appendix C.2 and assumes deterministic compressors, but the same guarantees hold under expectation with stochastic compressors, as is explained in Subsection˜C.3.
This second theoretical result analytically confirms the surprising numerical observation from Subsection˜3.1, illustrated in Figure˜1: $\mathrm{EF}$ and $\mathrm{EF^{21}}$ have the exact same optimal guarantee. Furthermore, the optimal step size is also the same, and can also be argued to be worst-case optimal in our setting using the same arguments as in Section 3.2. However, the optimal Lyapunov function is very different.
@@ 3.4 Tightness over multiple iterations and choice of state variables
A natural question arising from the above analysis concerns the tightness of the Lyapunov functions provided in Theorems˜1 and 2 over multiple steps. Specifically, for $K\geqslant 2$, we investigate whether the convergence rate of a method $\mathcal{M}\in{\mathrm{EF},\mathrm{EF^{21}},\mathrm{CGD}}$ can be improved by analyzing $\mathcal{M}^{K}$, the method run over $K$ iterations, defined as follows:
@@
We provide a numerical answer to that question in Appendix D.2 for all algorithms discussed in the paper, showing that the analysis of the single-state Lyapunov functions used in this work are tight even if we consider multiple iterations. To that end, we plot the worst-case contractions for multiple iterations computed using PEPit [58], on the optimal Lyapunov functions given by (11) and (13).
@@ 4 Methodology
We now present the methodology used to obtain the results in Section 3. The approach builds on the framework developed by [46]. We extend it to cover methods using deterministic compression operators under Assumption 1 in Appendix A. Obtaining the proofs of Theorems˜1 and 2 required a combination of advanced performance estimation techniques (finding optimal Lyapunov functions), several tricks, as well as symbolic computation and symbolic regression frameworks.
@@ 5 Conclusion
Our analysis is confined to the single-agent setting, both as an interesting problem on its own, and as a source of intuition for the multi-agent case. $\mathrm{CGD}$ cannot serve as a baseline in the multi-agent setting as it fails to converge with more than one agent [24]. In contrast, $\mathrm{EF^{21}}$ was specifically designed to improve convergence over $\mathrm{EF}$ in the multi-agent setting. Yet, its convergence rate in the single-agent case matches exactly that of $\mathrm{EF}$. The findings of this paper raise two compelling questions:
@@ Appendix A Feasibility problems with compressors
This section presents the methodology used to search for Lyapunov functions. In a nutshell, we formulate the Lyapunov search problem as a quasi-convex optimization problem involving linear matrix inequalities. Those problems are typically solved through the use of an iterative procedure involving a binary search with semidefinite solvers—we use MOSEK [57] throughout. The main steps taken here can be viewed as a generalization of the procedure proposed in [46] to first-order methods using compression, in particular Algorithms˜2 and 3. We also simplify a few steps that are not needed for our purposes. We start by reviewing the technique on a simpler example in Subsection˜A.1 before detailing the more tricky formulations involving compression.
@@ A.1 Feasibility problem for gradient descent
To introduce the concepts underlying the techniques used to construct Lyapunov functions for Algorithms˜2 and 3, we begin with the simpler case of gradient descent on smooth, strongly convex functions. That is, we consider the algorithm:
@@
$0\geqslant\inf_{\lambda_{ij}\geqslant 0}$ $0$ (26)
@@ A.2 Feasibility problem for $\mathrm{EF}$
Using the same ideas as in Subsection˜A.1, this section states the feasibility problem we solve to identify Lyapunov functions for Algorithm˜2. In short, this requires adapting the two steps presented in Subsection˜A.1 to accommodate the compressed message. Practically speaking, Step 1 must be adapted to a slightly larger problem (with a few more states both in the Gram matrix $G$ and in the Lyapunov candidates to incorporate compression). Then, Step 2 follows directly by the same reasoning as before: the condition derived in Step 1 reduces to checking the feasibility of a linear matrix inequality that is linear in $(P,p)$, which in turn allows us to search for the Lyapunov function via binary search over $\rho$.
@@
Consider running Algorithm˜2 with a deterministic compression operator satisfying Assumption˜1 for some $\epsilon\in[0,1]$ on any function satisfying Assumptions˜2 and 3. There exists a nonzero candidate Lyapunov function $\mathcal{V}_{(P,p)}$ of the form defined in (27), satisfying a given rate $\rho>0$, if and only if the following problem is feasible:
@@
The proof is decomposed into several steps that correspond to adapting the technical ingredients from Subsection˜A.1.
@@
$\Delta V_{P}(\rho)\coloneqq\begin{bmatrix}\bar{x}_{1}-\bar{x}_{\star}\\ \bar{g}_{1}-\bar{g}_{\star}\\ \bar{c}_{1}-\bar{c}_{\star}\\ \bar{e}_{1}-\bar{e}_{\star}\end{bmatrix}^{\top}P\begin{bmatrix}\bar{x}_{1}-\bar{x}_{\star}\\ \bar{g}_{1}-\bar{g}_{\star}\\ \bar{c}_{1}-\bar{c}_{\star}\\ \bar{e}_{1}-\bar{e}_{\star}\end{bmatrix}-\rho\begin{bmatrix}\bar{x}_{0}-\bar{x}_{\star}\\ \bar{g}_{0}-\bar{g}_{\star}\\ \bar{c}_{0}-\bar{c}_{\star}\\ \bar{e}_{0}-\bar{e}_{\star}\end{bmatrix}^{\top}P\begin{bmatrix}\bar{x}_{0}-\bar{x}_{\star}\\ \bar{g}_{0}-\bar{g}_{\star}\\ \bar{c}_{0}-\bar{c}_{\star}\\ \bar{e}_{0}-\bar{e}_{\star}\end{bmatrix}$$\quad$ (33)
@@
The interpolation conditions that enforce $f\in\mathcal{F}_{\mu,L}$ are identical to those we define in (24), but using the new basis vectors we define specifically for $\mathrm{EF}$. We do, however, need to introduce a new interpolation condition to encode the fact that we are using a contractive compressor. We do this for deterministic compression operators. Using our basis vectors, this corresponds to introducing the matrices
@@
Finally, following the same reasoning as described in Step 1 and Step 2 of Subsection˜A.1, and using the technical modifications we outlined here, we arrive at the feasibility problem described in the statement of the lemma. ∎
@@ A.3 Feasibility problem for $\mathrm{EF^{21}}$
As in the previous section, one can now adapt the same ideas as in Subsection˜A.1 and Subsection˜A.2 to $\mathrm{EF^{21}}$. This section states the feasibility problem we solve to identify Lyapunov functions in this context.
@@
Lemma 2 ($\mathrm{EF^{21}}$ feasibility problem).
Consider running Algorithm˜3 with a deterministic compression operator satisfying Assumption˜1 for some $\epsilon\in[0,1]$ on any function satisfying Assumptions˜2 and 3. There exists a nonzero candidate Lyapunov function $\mathcal{V}_{(P,p)}$ of the form defined in (38), satisfying a given rate $\rho>0$, if and only if the following problem is feasible:
@@
where matrices $(M_{ij})_{i,j\in\{0,1,\star\}}$ are defined as in Eq.˜24, $(C_{i}^{\mathrm{EF^{21}}})_{i\in\{0,1\}}$ given below in (44), and $\Delta V_{P}(\rho),\Delta v_{p}(\rho)$ are given below in (45) and (46). Here $\geqslant_{2}$ denotes coordinate-wise inequality in $\mathbb{R}^{2}$.
@@
Finally, using Step 1 and Step 2 of Subsection˜A.1, we arrive at the feasibility problem described in the statement of the lemma. ∎
Appendix B Rate comparison
In this section, we compare the rate given in equation˜14 with the convergence rate of $\mathrm{CGD}$ and the standard convergence rate of $\mathrm{EF^{21}}$ [14, see Theorem 2]. We show that $\mathrm{CGD}$ is strictly superior to $\mathrm{EF}$$\mathrm{EF^{21}}$ using a CAS certificate, that our rate is strictly better than the existing rate for $\mathrm{EF^{21}}$, and then we visualize the gap in both cases in Figures˜2 and 4.
We also compare the "complexity" resulting from either method in Figures˜3 and 5. In order to create a fair and interpretable comparison we used the following metric:
$\frac{\log\rho_{A}}{\log\rho_{B}}.$ (47)
The log-ratio can be motivated by the following example: if $\rho_{A}=0.99$ and $\rho_{B}=0.98$, then the latter will require half the number of iterations of the former before $\mathcal{V}(\xi_{N},x_{N};f)\leqslant\epsilon$.
B.1 Comparison between $\mathrm{CGD}$ and $\mathrm{EF}$$\mathrm{EF^{21}}$
We now compare the convergence rate when using the optimal step size between $\mathrm{CGD}$ and $\mathrm{EF}$ (equivalently, $\mathrm{EF^{21}}$ since the rate is the same). Note that the optimal rate for $\mathrm{CGD}$ is given by [64, 65]
$\rho_{\mathrm{CGD}}=\left(\frac{\kappa_{\sqrt{\epsilon}}-1}{\kappa_{\sqrt{\epsilon}}+1}\right)^{2},$ (48)
where $\kappa_{\sqrt{\epsilon}}=\kappa\left(\frac{1+\sqrt{\epsilon}}{1-\sqrt{\epsilon}}\right)$, when using step size [64]
$\eta_{\mathrm{CGD}}=\frac{2}{(1-\sqrt{\epsilon})\mu+(1+\sqrt{\epsilon})L}.$ (49)
The next statement has been verified using a WolframScript:
Certified using WolframScriptCertified using WolframScriptCAS  2 CAS  2 $\rho_{\mathrm{EF}}-\rho_{\mathrm{CGD}}=\sqrt{\epsilon}+$ $\left(\frac{1-\sqrt{\epsilon}}{2}\right)\left(\frac{\kappa-1}{\kappa+1}\right)^{2}\left[1-\sqrt{\epsilon}+\sqrt{(1+\sqrt{\epsilon})^{2}+\sqrt{\epsilon}16\frac{\kappa}{(\kappa-1)^{2}}}\right]$ - $\left(\frac{\kappa_{\sqrt{\epsilon}}-1}{\kappa_{\sqrt{\epsilon}}+1}\right)^{2}>0.$
B.2 Comparison to Richtárik et al. [14]
The rate of [14] for $\mathrm{EF^{21}}$ under the Łojasiewicz inequality, and with the largest possible step size chosen, is given by
$\rho^{\prime}\coloneqq\max\left\{1-\frac{1-\sqrt{\epsilon}}{2},\ 1-\kappa^{-1}\left(\frac{1-\sqrt{\epsilon}}{1+\sqrt{\epsilon}}\right)\right\}$ (50)
in the single-agent case. We begin by noting that to show that our rate is faster, we only have to show that it is smaller than one of the terms of equation˜50.
Certified using WolframScriptCertified using WolframScriptCAS  2 CAS  2 Note that $\rho^{\prime}-\rho$ $=1-\sqrt{\epsilon}-\frac{1-\sqrt{\epsilon}}{(1+\sqrt{\epsilon})\,\kappa}$ $-\frac{(1-\sqrt{\epsilon})(\kappa-1)^{2}\left(1-\sqrt{\epsilon}+\sqrt{\,1+\epsilon+\dfrac{2\sqrt{\epsilon}\,\bigl(1+\kappa(6+\kappa)\bigr)}{(\kappa-1)^{2}}}\,\right)}{2(\kappa+1)^{2}}$ Writing this as a single fraction, we see that the sign of this expression is the same as that of $\Delta\coloneqq\bigl(1-\sqrt{\epsilon}\bigr)\Bigl(-2+\kappa\bigl(-3+\epsilon(\kappa-1)^{2}+2\sqrt{\epsilon}(\kappa+1)^{2}+(1+\sqrt{\epsilon})R$ $\qquad\qquad\quad+\kappa\bigl(4+\kappa-(1+\sqrt{\epsilon})R\bigr)\bigr)\Bigr).$ where $R\coloneqq\sqrt{\,(\kappa-1)^{2}+\epsilon(\kappa-1)^{2}+2\sqrt{\epsilon}\,\bigl(1+\kappa(6+\kappa)\bigr)}$. Denote the coefficient in front of $R$ as $C(R)$, and note that is negative. As a result, it just remains to make sure that the rest of the terms compensate for this term. Rewriting this as an inequality, and squaring to deal with $R$, our result follows from the inequality $\left(\Delta-C(R)\cdot R\right)^{2}-\left(C(R)\cdot R\right)^{2}=4(1-\sqrt{\epsilon})^{2}(1+\kappa)F>0,$ where $F=1+\kappa\left(1+\kappa(-5+3\kappa)+\epsilon(1+\kappa)(-1+3\kappa)+2\sqrt{\epsilon}\,(-1+\kappa)(1+3\kappa)\right).$ The factors other than $F$ are clearly positive, and F is increasing with respect to $\epsilon$. Setting $\epsilon$ to 0 shows that it is positive for all valid $\epsilon$.
Figure 2: Line plot comparing the convergence rate of this paper (blue) with equation˜50 (red) for various $\kappa$.
Figure 3: Line plot comparing the complexity of equation˜50 with the rates of this paper for various $\kappa$.
Figure 4: Line plot comparing the convergence rate of this $\mathrm{CGD}$ (blue) with $\mathrm{EF}$$\mathrm{EF^{21}}$ (red) for various $\kappa$.
Figure 5: Line plot comparing the complexity of $\mathrm{EF}$$\mathrm{EF^{21}}$ with $\mathrm{CGD}$ for various $\kappa$.
Appendix C Missing proofs